Tom is 8. His mom is 31. How long before she is twice his age?
Answer by kapilsinghi(68)

Tom is 8. His mom is 31. How long before she is twice his age?
let x years from now age of mother is twice the age of tom
so 2*(8+x) = 31+x
16+2x = 31 +x
x= 15
kapilsinghi123@gmail.com